### v5

@@ Abstract
Modern theories of phase transitions and scale-invariance are rooted in path integral formulation and renormalization groups (RG). Despite the applicability of these approaches on simple systems with only pairwise interactions, they are less effective on complex systems with un-decomposable high-order interactions (i.e., interactions among arbitrary sets of units). To precisely characterize the universality of high-order interacting systems, we propose simplex path integral and simplex renormalization group (SRG) as the generalizations of classic approaches to arbitrary high-order and heterogeneous interactions. We first formalize the trajectories of units governed by high-order interactions to define path integrals on corresponding simplexes based on a high-order propagator. Then we develop a method to integrate out short-range high-order interactions in the momentum space, accompanied by a coarse graining procedure functioning on the simplex structure generated by high-order interactions. The proposed SRG, equipped with a divide-and-conquer framework, can deal with the absence of ergodicity arised from the sparse distribution of high-order interactions and renormalize a system with intertwined high-order interactions on $p$-order according to its properties on $q$-order ($p\leq q$). The associated scaling relation and its corollaries support to differentiate among scale-invariant, weakly scale-invariant, and scale-dependent systems across different orders. We have validated our theory in multi-order scale-invariance verification, topological invariance discovery, organizational structure identification, and information bottleneck analysis to demonstrate its capacity for identifying intrinsic properties of high-order interacting systems during system reduction.
@@ I.1 Unknowns about high-order interactions
However, a non-negligible theoretical vacancy can be found in the existing path integral and renormalization group approaches if we subdivide interactions into pairwise and high-order categories. As the name suggests, a pairwise interaction only involves a pair of units and does not require the participation of any other unit. This kind of interaction can be represented by edges in networks [8, 9] and are implicitly used in the derivations of classic path integrals and renormalization groups (e.g., see Ref. [10]). A high-order interaction, on the contrary, is the mutual coupling among more than two units [11, 12, 13]. While some high-order interactions are reducible and can be decomposed into a group of pairwise interactions, most high-order interactions are un-decomposable and in-equivalent to the direct sum of the pairwise ones [9, 14]. For instance, the triplet collaboration among three agents is not equivalent to the trivial sum of three pairs of individual collaborations. These un-decomposable high-order interactions have intricate effects on system dynamics [13, 14] and can not be trivially represented by ordinary networks [12]. Although notable efforts have been devoted to studying the optimal characterization of high-order interactions (e.g., using simplicial complexes [15, 16, 17, 18] or hypergraphs [19, 20, 21]), there are fewer works focusing on developing specialized path integral and renormalization group theories for complex systems with high-order interactions [22, 17]. The difficulty of proposing these specialized frameworks arises from the intrinsic dependence on pairwise interactions when researchers constrain the discrete replicas of interacting systems as conventional networks. Once this constraint is relaxed, the intertwined effects of newly included high-order interactions would make the analysis on emerging dynamics highly non-trivial and even lead to novel results that can not be directly derived by classic theories (e.g., see instances in diffusion and random walks [16, 23, 20], social dynamics [24, 25, 26], and neural dynamics [27, 28]).
@@ I.2 Related works and remaining challenges
To date, most progress of the computational implementations of path integrals and renormalization groups is achieved on real systems with pairwise interactions [29, 30, 10, 31, 32, 33]. Although box-covering methods [34, 35, 36, 37] may be the most straightforward ways for coarse graining while preserving the organizational properties of a system represented by a network, they depend on the assumption of internal fractal properties and convey no information about system dynamics. Different from box-covering, spectral method [38] pays more attention to inherent dynamics properties (e.g., random walks) during coarse graining but may suffer from the obstacles of finding super-units implied by small-world effects [39, 33, 10]. Aiming at resolving the coexisting and correlated scales, a geometric renormalization group is developed to identify the potential geometric scaling rather than path-distance scaling [33]. Nevertheless, this approach essentially relies on the a-priori-knowledge of the hidden metric space of networks (e.g., the hyperbolic space [40]). With the pursuit of realizing the metric-free coarse graining whose iterability is not limited by small-world effects, a Laplacian renormalization group (LRG) [10] has been derived from the diffusion on networks, a continuum counterpart of the block-spin transformation [41], to deal with the network topology, dynamics, and geometry encoded by the graph Laplacian operator [42]. This approach has a natural connection with path integrals as well [43]. Although these properties make the LRG a promising foundation for in-depth explorations, this framework suffers from several non-negligible limitations. First, there is no apparent correspondence between the LRG and an appropriate approach for analyzing un-decomposable high-order interactions because the Laplacian operator can not characterize polyadic relations directly [11, 12, 13, 9, 14]. Second, the intrinsic dependence of the LRG on the ergodic property of system dynamics reflected by network connectivity limits the applicability of this approach in analyzing high-order interactions, which are more sparsely distributed in the system than pairwise interactions and do not necessarily ensure ergodicity.
@@
There exist fewer pioneering works devoted to developing renormalization groups for high-order interactions, among which, a notable framework is the real space renormalization group proposed by Refs. [22, 17]. This approach is initially defined on the Laplacian operators of the network skeletons of the Apollonian [44] and the pseudo-fractal [45] simplicial complexes to calculate spectral dimensions [22]. Then, it is generalized to the normalized up-Laplacian operator [23, 16] of simplicial complexes to deal with high-order spectra. Despite the effectiveness of this approach, its dependence on the non-trivial manual derivations using the Gaussian model [46] and the specialized application scope about spectral dimensions [22, 17] make it less applicable to the real scenarios where a programmatic implementation is demanded for renormalizing the complex systems generated by empirical data automatically.
@@ I.3 Our frameworks and contributions
Our work suggests a way to define these concepts on simplexes, in which simplexes can join any sub-set of units to characterize un-decomposable high-order interactions among them (see Fig. 1(a) for instances and Appendix A for definitions). We propose two kinds of high-order propagators as the generalized path integral formulation of the abstract diffusion processes driven by multi-order interactions, which further creates a system mode description tantamount to the moment space. The coarse graining in the moment space is implemented by progressively integrating out short-range system modes corresponding to some certain order of interactions. Parallel to this process, we propose a real space coarse graining procedure to reduce the simplex structures associated with short-range high-order interactions, where we suggest a way to overcome the dependence on ergodicity. The correspondence between the moment and real spaces enables our framework to the system coarse graining according to both organizational structure and internal dynamics. Taken together, these definitions give birth to our simplex path integral and simplex renormalization group (SRG) theories, which are applicable to arbitrary orders of interactions.
@@
To comprehensively demonstrate the capacity of our approaches in identifying the latent characteristics of diverse real complex systems, we validate our framework from the perspectives of multi-order scale-invariance detection, topological invariance discovery, latent organizational structure identification, and information bottleneck optimization, respectively. An efficient code implementation of this framework is provided in [51]. One can also see a detailed explanation about the programmatic implementation and usage of the SRG in Appendix K.
@@ II Un-decomposable high-order interactions
The Laplacian operator is a natural choice for characterizing interactions and their spreading among units [52, 53, 54]. In a special case with pairwise interactions, the Laplacian has a classic expression, $\mathbf{L}_{ij}=\delta_{ij}\sum_{k}\mathbf{A}_{ik}-\mathbf{A}_{ij}$, where $\delta$ denotes the Kronecker delta function and $\mathbf{A}$ is a weighted adjacency matrix that describes the non-negative coupling strengths among all units in unit set $V$ (here the non-negativity is required by the Laplacian). Under the autonomous condition, the evolution of system dynamics given an initial state $\mathbf{x}_{0}$ is defined by $\mathbf{x}=\exp\left(-\tau\mathbf{L}\right)\mathbf{x}_{0}$, where $\tau\in\left(0,\infty\right)$ denotes a time scale [55]. However, in more general cases with un-decomposable high-order interactions, the classic Laplacian operator is no longer applicable. To fill the gap, previous studies have explored the generalization of the Laplacian operator on uniform hypergraphs [56, 57] and simplicial complexes [58, 59, 23, 9], where diverse variants of the Laplacian (e.g., the combinatorial Laplacian [60], the Hodge Laplacian [61, 23], and the multi-order Laplacian [9]) have been proposed to characterize high-order interactions. In our work, we first derive a multi-order Laplacian following the spirit of Ref. [9]. Then we develop a new operator, referred to as the high-order path Laplacian, as an alternative description of high-order interactions from a different perspective. As shown in Fig. 1, these two operators are defined on clique complexes, a kind of commonly used simplicial complexes in the persistent homology analysis [62, 60] (see Appendix A).
@@ II.1 High-order interactions through faces
Let us consider the $n$-order interactions that manifest as the simultaneous participation of $n+1$ related units. The un-decomposability of these interactions lies in that the couplings among $n+1$ units are required to occur simultaneously and none of them is dispensable. For instance, the triplet collaboration requires the participation of all three agents together. As discussed in Ref. [9], these interactions can be represented by the $n$-faces of the $n$-simplex where $n+1$ units are placed on. For the sake of clarity, we define $\{i\}_{n+1}\subset V$ as an arbitrary set of $n+1$ units in which unit $i$ is included and denote $S^{\{i\}}_{n+1}$ as the set of all permutations (i.e., rearrangements) on $\{i\}_{n+1}$. The multi-order Laplacian $\mathbf{L}_{M}^{\left(n\right)}$ associated with $n$-simplexes is defined as
@@
In Eq. (1), term $\mathbf{D}_{i}^{\left(n\right)}$ counts the number of $n$-simplexes that contains unit $i$
$\mathbf{D}_{i}^{\left(n\right)}=\frac{1}{\left(n+1\right)!}\sum_{\{i\}_{n+1}\subset V}\prod_{\omega\in S^{\{i\}}_{n+1}}\mathbf{M}_{\omega},$ (2)
where $\frac{1}{\left(n+1\right)!}$ serves the normalization for high-order simplexes (i.e., the enumeration in Eq. (2) involves repetition in high-order cases). Term $\mathbf{A}_{ij}^{\left(n\right)}$ counts the number of $n$-simplexes that contains units $i$ and $j$
$\mathbf{A}_{ij}^{\left(n\right)}=\frac{1-\delta_{ij}}{\left(n+1\right)!}\sum_{\{i,j\}_{n+1}\subset V}\prod_{\omega\in S^{\{i,j\}}_{n+1}}\mathbf{M}_{\omega},$ (3)
@@
where $\mathbf{A}=\mathbf{A}^{\left(1\right)}$ is the classic adjacency matrix of units.
In Eq. (2), vector $\mathbf{D}^{\left(n\right)}$ is similar to the degree vector while matrix $\mathbf{A}^{\left(n\right)}$ can be understood as the adjacency matrix. In a special case where $n=1$, Eq. (2) is equivalent to the classic Laplacian for pairwise interactions. In the more general cases where $n>1$, high-order interactions can be characterized by Eq. (2). Please see Fig. 1(d) for the illustrations of the multi-order Laplacian. Although the definitions of Eqs. (2-4) are different from the approach proposed by Ref. [9], they are mathematically consistent with Ref. [9].
@@ II.2 High-order interactions along paths
Then, we turn to another type of $n$-order interactions, which manifest as the sequential actions of $n+1$ related units after they all participate in. The un-decomposability of these interactions arises from that the mutual coupling effects emerge only after $n+1$ units have been engaged with the interplay progressively. For example, a pipeline-like collaboration may require all agents to communicate with each other and behave in order. Each previous action is globally known to all agents and affects all subsequent actions. This property can be characterized by the paths (i.e., a sequence of $1$-simplexes) that successively pass through the $n+1$ units placed on the $n$-simplex. To describe the interactions propagating along the paths on $n$-simplexes, we develop an operator named as the high-order path Laplacian $\mathbf{L}_{H}^{\left(n\right)}$
@@
In Eq. (5), coefficient $\frac{1}{n}$ denotes the normalization of the propagation speed of interactions along paths. Term $\mathbf{P}_{i}^{\left(n\right)}$ counts the number of paths that traverse all $n+1$ units on the $n$-simplex without repetition and terminate at unit $i$
$\mathbf{P}_{i}^{\left(n\right)}=\frac{1}{n+1}\sum_{\{i\}_{n+1}\subset V}\prod_{\omega\in S^{\{i\}}_{n+1}}\mathbf{M}_{\omega},$ (6)
@@
where $\frac{1}{n+1}$ denotes the normalization. Term $\mathbf{B}_{ij}^{\left(n\right)}$ counts the number of paths that traverse all $n+1$ units on the $n$-simplex without repetition, initiate at unit $j$, and terminate at unit $i$
$\mathbf{B}_{ij}^{\left(n\right)}=\frac{1-\delta_{ij}}{n\left(n+1\right)}\sum_{\{i,j\}_{n+1}\subset V}\prod_{\omega\in S^{\{i,j\}}_{n+1}}\mathbf{M}_{\omega}.$ (7)
@@
Please see Fig. 1(e) for the instances of the high-order path Laplacian. In the case with $n=1$, Eqs. (6-7) reduce to the classic degree vector and adjacency matrix, respectively (i.e., we have $\mathbf{P}^{\left(1\right)}=\mathbf{D}^{\left(1\right)}$ and $\mathbf{B}^{\left(1\right)}=\mathbf{A}$). Therefore, Eq. (5) denotes the classic Laplacian when interactions are pairwise.
@@
In general, one can choose the multi-order Laplacian $\mathbf{L}_{M}^{\left(n\right)}$ and the high-order path Laplacian $\mathbf{L}_{H}^{\left(n\right)}$ according to application demands. While $\mathbf{L}_{M}^{\left(n\right)}$ conveys information about the high-order interactions that units are simultaneously engaged with, the definition of $\mathbf{L}_{H}^{\left(n\right)}$ is more applicable to the cases where a high-order interaction is formed by an irreducible sequence of the participation of related units. We offer efficient code implementations of these two kinds of operators, which can be seen in [51]. Meanwhile, in Appendix B, we explain the similarities and differences between our considered Laplacian operators and other Laplacian operators used in topology theories (e.g., the combinatorial Laplacian [60] and the Hodge Laplacian [61, 23] in persistent homology analysis [63, 64, 62, 65, 66, 67]), which helps us understand the theoretical significance of designing renormalization groups on $\mathbf{L}_{M}^{\left(n\right)}$ and $\mathbf{L}_{H}^{\left(n\right)}$. For convenience, we uniformly denote them as $\mathbf{L}^{\left(n\right)}$ in our subsequent derivations. One can specify the actual definition of $\mathbf{L}^{\left(n\right)}$ in the application.
@@
Figure 2: Illustrations of the simplex path integral. (a) An interacting system of $100$ units is generated, where the pairwise interactions of units follow a Barabási-Albert network ($c=6$). This system is used to show the derivation of the time evolution operators of the first three orders, where we use the multi-order Laplacian (MOL) for calculation and set $\tau=1$. For each order of interactions, we illustrate the first four terms of the series expansion of time evolution operator in Eq. (5). (b) For the system defined in (a), we show the density operators calculated based on the MOL or the high-order path Laplacian (HOPL) on each order. (c) We show that values in the density operator can be related to the evolution paths of system states. The blue path leads from one state to another while the red path starts from and ends at the same state.

### v7

@@ Abstract
Modern theories of phase transitions and scale-invariance are rooted in path integral formulation and renormalization group (RG). Despite the applicability of these approaches on simple systems with only pairwise interactions, they are less effective on complex systems with un-decomposable high-order interactions (i.e., interactions among arbitrary sets of units). To precisely characterize the universality of high-order interacting systems, we propose simplex path integral and simplex renormalization group (SRG) as the generalizations of classic approaches to arbitrary high-order and heterogeneous interactions. We first formalize the trajectories of units governed by high-order interactions to define path integrals on corresponding simplices based on a high-order propagator. Then we develop a method to integrate out short-range high-order interactions in the momentum space, accompanied by a coarse graining procedure functioning on the simplex structure generated by high-order interactions. The proposed SRG, equipped with a divide-and-conquer framework, can deal with the absence of ergodicity arised from the sparse distribution of high-order interactions and renormalize a system with intertwined high-order interactions on the $p$-order according to its properties on the $q$-order ($p\leq q$). The associated scaling relation and its corollaries support to differentiate among scale-invariant, weakly scale-invariant, and scale-dependent systems across different orders. We have validated our theory in multi-order scale-invariance verification, topological invariance discovery, organizational structure identification, and information bottleneck analysis. These experiments demonstrate the capacity of our theory for identifying intrinsic statistical and topological properties of high-order interacting systems during system reduction.
@@ I.1 Unknowns about high-order interactions
However, a theoretical vacancy can be found in the existing path integral and renormalization group approaches if we subdivide interactions into pairwise and high-order categories. As the name suggests, a pairwise interaction only involves a pair of units and does not require the participation of any other unit. This kind of interactions can be represented by edges in networks [8, 9] and are implicitly used in the derivations of classic path integral and renormalization group theories (e.g., see Ref. [10]). A high-order interaction, on the contrary, is the mutual coupling among more than two units [11, 12, 13]. While some high-order interactions can be decomposed into a group of pairwise interactions, most high-order interactions are un-decomposable and in-equivalent to the direct sum of the pairwise ones [9, 14]. For instance, the triplet collaboration among three agents is not equivalent to the trivial sum of three pairs of individual collaborations. These un-decomposable high-order interactions have intricate effects on system dynamics [13, 14] and can not be trivially represented by ordinary networks [12]. Although notable efforts have been devoted to studying the optimal characterization of high-order interactions (e.g., using simplicial complexes [15, 16, 17, 18] or hypergraphs [19, 20, 21]), there are fewer works focusing on developing specialized path integral and renormalization group theories for complex systems with high-order interactions [22, 17]. The difficulty of proposing these specialized frameworks arises from the intrinsic dependence on pairwise interactions when researchers constrain the discrete replicas of interacting systems as conventional networks. Once this constraint is relaxed, the intertwined effects of newly included high-order interactions would make the analysis on emerging dynamics highly non-trivial and even lead to novel results that can not be directly derived by classic theories (e.g., see instances in diffusion and random walks [16, 23, 20], social dynamics [24, 25, 26], and neural dynamics [27, 28]).
@@ I.2 Related works and remaining challenges
To date, most progress of the computational implementations of path integral and renormalization group is achieved on real systems with pairwise interactions [29, 30, 10, 31, 32, 33]. Although box-covering methods [34, 35, 36, 37] may be the most straightforward ways for coarse graining while preserving the organizational properties of a system represented by a network, they depend on the assumption of internal fractal properties and convey no information about system dynamics. Different from box-covering, spectral method [38] pays more attention to inherent dynamics properties (e.g., random walks) during coarse graining but may suffer from the obstacles of finding macro-units implied by small-world effects [39, 33, 10]. Aiming at resolving the coexisting and correlated scales, a geometric renormalization group is developed to identify the potential geometric scaling rather than path-distance scaling [33]. Nevertheless, this approach essentially relies on the a-priori-knowledge of the hidden metric space of networks (e.g., the hyperbolic space [40]). With the pursuit of realizing the metric-free coarse graining whose iterability is not limited by small-world effects, a Laplacian renormalization group (LRG) [10] has been derived from the diffusion on networks, a continuum counterpart of the block-spin transformation [41], to deal with the network topology, dynamics, and geometry encoded by the graph Laplacian operator [42]. This approach has a natural connection with path integrals as well [43]. Although these properties make the LRG a promising foundation for in-depth explorations, this framework suffers from several non-negligible limitations. First, there is no apparent correspondence between the LRG and an appropriate approach for analyzing un-decomposable high-order interactions because the Laplacian operator can not characterize polyadic relations directly [11, 12, 13, 9, 14]. Second, the intrinsic dependence of the LRG on the ergodic property of system dynamics reflected by network connectivity limits the applicability of this approach in analyzing high-order interactions, which are more sparsely distributed in the system than pairwise interactions and do not necessarily ensure ergodicity.
@@
There exist fewer pioneering works devoted to developing renormalization group theories for high-order interactions, among which, a notable framework is the real space renormalization group proposed by Refs. [22, 17]. This approach is initially defined on the Laplacian operators of the network skeletons of the Apollonian [44] and the pseudo-fractal [45] simplicial complexes to calculate spectral dimensions [22]. Then, it is generalized to the normalized up-Laplacian operator [23, 16] of simplicial complexes to deal with high-order spectra. Despite the effectiveness of this approach, its dependence on the non-trivial manual derivations using the Gaussian model [46] and the specialized application scope about spectral dimensions [22, 17] make it less applicable to the real scenarios where a programmatic implementation is demanded for renormalizing the complex systems generated by empirical data automatically.
@@ I.3 Our frameworks and contributions
Our work suggests a way to define these concepts on simplices, in which simplices can join any sub-set of units to characterize un-decomposable high-order interactions among them (see Fig. 1(a) for instances and Appendix A for definitions). We propose two kinds of high-order propagators as the generalized path integral formulation of the abstract diffusion processes driven by multi-order interactions, which further creates a system mode description tantamount to the moment space. The coarse graining in moment space is implemented by progressively integrating out short-range system modes corresponding to some certain order of interactions. Parallel to this process, we propose a real space coarse graining procedure to reduce the simplex structures associated with short-range high-order interactions, where we suggest a way to overcome the dependence on ergodicity. The correspondence between the moment and real spaces enables our framework to the system coarse graining according to both organizational structure and internal dynamics. Taken together, these definitions give birth to our simplex path integral and simplex renormalization group (SRG) theories, which are applicable to arbitrary orders of interactions.
@@
To comprehensively demonstrate the capacity of our approaches in identifying the latent characteristics of diverse real complex systems, we validate our framework from the perspectives of multi-order scale-invariance detection, topological invariance discovery, latent organizational structure identification, and information bottleneck optimization, respectively. An efficient code implementation of this framework is provided in [51]. One can also see a detailed explanation about the programmatic implementation and usage of the SRG in Appendix M.
@@ II Un-decomposable high-order interactions
The Laplacian operator is a natural choice for characterizing interactions and their spreading among units [52, 53, 54]. In a special case with pairwise interactions, the Laplacian has a classic expression, $\mathbf{L}_{ij}=\delta_{ij}\sum_{k}\mathbf{A}_{ik}-\mathbf{A}_{ij}$, where $\delta$ denotes the Kronecker delta function and $\mathbf{A}$ is a weighted adjacency matrix that describes the non-negative coupling strengths among all units in unit set $V$ (here the non-negativity is required by the Laplacian). Under the autonomous condition, the evolution of system dynamics given an initial state $\mathbf{x}_{0}$ is defined by $\mathbf{x}=\exp\left(-\tau\mathbf{L}\right)\mathbf{x}_{0}$, where $\tau\in\left(0,\infty\right)$ denotes a time scale [55]. However, in more general cases with un-decomposable high-order interactions, the classic Laplacian operator is no longer applicable. To fill the gap, previous studies have explored the generalization of the Laplacian operator on uniform hypergraphs [56, 57] and simplicial complexes [58, 59, 23, 9], where diverse variants of the Laplacian (e.g., the combinatorial Laplacian [60], the Hodge Laplacian [61, 23], and the multi-order Laplacian [9]) have been proposed to characterize high-order interactions.
In our work, we first derive a multi-order Laplacian following the spirit of Ref. [9]. Then, we develop a new operator, referred to as the high-order path Laplacian, as an alternative description of high-order interactions from a different perspective. As shown in Fig. 1, these two operators are defined on clique complexes, a kind of commonly used simplicial complexes in the persistent homology analysis [62, 60] (see Appendix A for clique complexes). The derivations of these two operators are elaborated in Appendix B, where we summarize the similarities and differences between them. Although these two operators have different physics meanings and derivation processes, they can be reformulated into similar mathematical forms as presented in Secs. II.1-II.2. Thus, the theoretical differences between these two operators will not impede their unified numerical calculations and programmatic implementations.
@@ II.1 High-order interactions through faces
Let us consider the $n$-order interactions that manifest as the simultaneous participation of $n+1$ related units. The un-decomposability of these interactions lies in that the couplings among $n+1$ units are required to occur simultaneously and none of them is dispensable. For instance, the triplet collaboration requires the participation of all three agents together. As discussed in Ref. [9], these interactions can be represented by the $n$-faces of the $n$-simplex where $n+1$ units are placed on. For the sake of clarity, we define $\{i\}_{n+1}\subset V$ as an arbitrary set of $n+1$ units in which unit $i$ is included and denote $S^{\{i\}}_{n+1}$ as the set of all permutations (i.e., rearrangements) on $\{i\}_{n+1}$. The multi-order Laplacian $\mathbf{L}_{M}^{\left(n\right)}$ associated with $n$-simplices is defined as
@@
In Eq. (1), the term $\mathbf{D}_{i}^{\left(n\right)}$ counts the number of $n$-simplices that contains unit $i$
$\mathbf{D}_{i}^{\left(n\right)}=\sum_{\{i\}_{n+1}\subset V}\prod_{\omega\in S^{\{i\}}_{n+1}}\mathbf{M}_{\omega},$ (2)
Meanwhile, the term $\mathbf{A}_{ij}^{\left(n\right)}$ counts the number of $n$-simplices that contains units $i$ and $j$
$\mathbf{A}_{ij}^{\left(n\right)}=\left(1-\delta_{ij}\right)\sum_{\{i,j\}_{n+1}\subset V}\prod_{\omega\in S^{\{i,j\}}_{n+1}}\mathbf{M}_{\omega},$ (3)
@@
where $\mathbf{A}=\mathbf{A}^{\left(1\right)}$ is the classic adjacency matrix of units. It can be seen that $\prod_{\omega\in S^{\{i,j\}}_{n+1}}\mathbf{M}_{\omega}$ serves as an indicator function, which equals $1$ if the units in $\{i,j\}_{n+1}$ form a $n$-simplex and equals 00 otherwise.
Please see Appendix B for the full derivations of the multi-order Laplacian. In Eq. (2), vector $\mathbf{D}^{\left(n\right)}$ is similar to the degree vector while matrix $\mathbf{A}^{\left(n\right)}$ can be understood as the adjacency matrix. The coefficient $n$ in Eq. (2) arises from the fact that $\sum_{j}\mathbf{A}_{ij}^{\left(n\right)}=n\mathbf{D}^{\left(n\right)}_{i}$ (i.e., each simplex is repeatedly counted $n$ times). In a special case where $n=1$, Eq. (2) is equivalent to the classic Laplacian for pairwise interactions. In the more general cases where $n>1$, high-order interactions can be characterized by Eq. (2). Please see Fig. 1(d) for the illustrations of the multi-order Laplacian. Although the definitions of Eqs. (2-4) are different from the approach proposed by Ref. [9], they are mathematically consistent with Ref. [9].
@@ II.2 High-order interactions along paths
Then, we turn to another type of $n$-order interactions, which manifest as the sequential actions of $n+1$ related units after they all participate in (i.e., sequential participation). The un-decomposability of these interactions arises from that the mutual coupling effects emerge only after $n+1$ units have been engaged with the interplay progressively. For example, a pipeline-like collaboration may require all agents to communicate with each other and behave in order. Each previous action is globally known to all agents and affects all subsequent actions. This property can be characterized by the paths (i.e., a sequence of $1$-simplices) that successively pass through the $n+1$ units placed on the $n$-simplex. Consequently, to model the $n$-order interactions arisen from sequential participation, we need to analyze both the faces of simplices (i.e., to ensure that all related units are indispensable) and the paths defined on these faces (i.e., to describe the sequential participation).
To describe the interactions propagating along the paths on $n$-simplices, we develop an operator named as the high-order path Laplacian $\mathbf{L}_{H}^{\left(n\right)}$
@@
In Eq. (5), the coefficient $\frac{1}{n}$ denotes the normalization of the propagation speed of interactions along paths. The term $\mathbf{P}_{i}^{\left(n\right)}$ counts the number of paths that traverse all $n+1$ units on the $n$-simplex without repetition and terminate at unit $i$
$\mathbf{P}_{i}^{\left(n\right)}=n!\sum_{\{i\}_{n+1}\subset V}\prod_{\omega\in S^{\{i\}}_{n+1}}\mathbf{M}_{\omega},$ (6)
@@
where $n!$ counts the possibilities for the rest part of units in $\{i\}_{n+1}$ to form unique sequences after we fix unit $v_{i}$ as the endpoint of these sequences. The term $\mathbf{B}_{ij}^{\left(n\right)}$ counts the number of paths that traverse all $n+1$ units on the $n$-simplex without repetition, initiate at unit $j$, and terminate at unit $i$
$\mathbf{B}_{ij}^{\left(n\right)}=\left(1-\delta_{ij}\right)\left(n-1\right)!\sum_{\{i,j\}_{n+1}\subset V}\prod_{\omega\in S^{\{i,j\}}_{n+1}}\mathbf{M}_{\omega},$ (7)
where the coefficient $\left(n-1\right)!$ is derived from the fact that the rest part of units in $\{i,j\}_{n+1}$ can form $\left(n-1\right)!$ unique sequences if we fix units $v_{i}$ and $v_{j}$ as the endpoints.
@@
The construction process of the high-order path Laplacian $\mathbf{L}_{H}^{\left(n\right)}$ is elaborated in Appendix B. Please see Fig. 1(e) for several instances of the high-order path Laplacian. In a special case with $n=1$, Eqs. (6-7) reduce to the classic degree vector and adjacency matrix, respectively (i.e., we have $\mathbf{P}^{\left(1\right)}=\mathbf{D}^{\left(1\right)}$ and $\mathbf{B}^{\left(1\right)}=\mathbf{A}$). Therefore, Eq. (5) denotes the classic Laplacian when interactions are pairwise. In most general cases with $n>1$, the $n$-order interactions formed by sequential participation are not equivalent to those caused by simultaneous participation (see Appendix B for details).
In general, one can choose the multi-order Laplacian $\mathbf{L}_{M}^{\left(n\right)}$ and the high-order path Laplacian $\mathbf{L}_{H}^{\left(n\right)}$ according to application demands. As we have suggested above, while $\mathbf{L}_{M}^{\left(n\right)}$ conveys information about the high-order interactions that units are simultaneously engaged with, the definition of $\mathbf{L}_{H}^{\left(n\right)}$ is more applicable to the cases where a high-order interaction is formed by an irreducible sequence of the participation of related units. In Appendix B, we explain how these two operators differ from each other in governing the dynamics of system evolution (e.g., with different decay rates on different orders). Specifically, the decay rate of system evolution maintains the same no matter units follow the simultaneous (i.e., $\mathbf{L}_{M}^{\left(n\right)}$) or the sequential (i.e., $\mathbf{L}_{H}^{\left(n\right)}$) participation when interactions are pairwise (i.e., $n=1$). When units exhibit $2$-order interactions, the sequential participation implies slower decays of system evolution compared with the simultaneous participation. When interactions happen on higher orders (i.e., $n>2$), the sequential participation leads to a larger decay rate. These differences highlight the dissimilar physics meanings and characteristics of $\mathbf{L}_{M}^{\left(n\right)}$ and $\mathbf{L}_{H}^{\left(n\right)}$.
@@
Based on the unified mathematical forms presented in Secs. II.1-II.2, we have developed efficient code implementations of these two kinds of operators, which can be seen in [51]. Meanwhile, in Appendix C, we explain the similarities and differences between our considered Laplacian operators and other Laplacian operators used in topology theories (e.g., the combinatorial Laplacian [60] and the Hodge Laplacian [61, 23] in persistent homology analysis [63, 64, 62, 65, 66, 67]), which helps us understand the theoretical significance of designing a renormalization group on $\mathbf{L}_{M}^{\left(n\right)}$ and $\mathbf{L}_{H}^{\left(n\right)}$. For convenience, we uniformly denote them as $\mathbf{L}^{\left(n\right)}$ in our subsequent derivations. One can specify the actual definition of $\mathbf{L}^{\left(n\right)}$ in the application.
@@
Figure 2: Illustrations of the simplex path integral. (a) An interacting system of $50$ units is generated, where the pairwise interactions of units follow a Barabási-Albert network ($c=4$). This system is used to show the derivation of the time evolution operators of the first three orders, where we use the multi-order Laplacian (MOL) for calculation and set $\tau=1$. For each order of interactions, we illustrate the first four terms of the series expansion of time evolution operator in Eq. (9). (b) For the system defined in (a), we show the density operators calculated based on the MOL or the high-order path Laplacian (HOPL) on each order. (c) We show that values in the density operator can be related to the evolution paths of system states. The blue path leads from one state to another while the red path starts from and ends at the same state.